• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 15:05:25 on July 15, 2006

    Entry number 174203

    Followups:

    keyword=Noisy wire left VDC

    First runs after the left VDC power supplies was fixed and timing was ok show a couple of noisy wires : U2 wires 35 to 39 U3 V1 wires 239 to 286 V15-V17

    looking at the cratemap, this should correspond to ROC3 slot 20 chan 0-47 ROC3 slot 6 chan chan 32-47

    I'll check the ribbon cable for those channels first ans signal with the oscilloscope.

    A copy of this log entry has been emailed to: bogdanw@jlab.org,segal@jlab.org



    Figure 1